OpenStudy (anonymous):

use logarithmic differntiation:

OpenStudy (anonymous):

to find y given : \[y=(\tan \theta)\sqrt{2\theta+1)}\]

OpenStudy (anonymous):

all you need to do is log both sides

OpenStudy (anonymous):

can you show?

OpenStudy (anonymous):

log y = logt(theta) + 1/2 log (2theta +1)

OpenStudy (anonymous):

im here to learn!

myininaya (myininaya):

\[ln(y)=\ln[\tan(\theta)*\sqrt{2\theta+1}]\] \[\ln(y)=\ln(\tan(\theta))+\ln(2\theta+1)^\frac{1}{2}\] \[\ln(y)=\ln(\tan(\theta))+\frac{1}{2}\ln(2\theta+1)\] take derivative of both sides \[\frac{y'}{y}=\frac{\sec^2(\theta)}{\tan(\theta)}+\frac{1}{2}*\frac{2}{2\theta+1}\]

OpenStudy (anonymous):

then when you diff log y you get 1/y *dy/dx =rhs

myininaya (myininaya):

now our objective is to find y' so multiply y on both sides and woolah hey joe did you learn something lol

OpenStudy (anonymous):

interesting! i have never heard of that technique before, but it makes sense lol

myininaya (myininaya):

lol i dont believe you

OpenStudy (anonymous):

totally serious! i was like, "logarithmic differentiation......wut"

myininaya (myininaya):

logarithmic differentiation can be really useful for some functions

OpenStudy (anonymous):

but its pretty much what it sounds like.

OpenStudy (anonymous):

can someone write it out for me:) i will be eternally grateful

OpenStudy (anonymous):

yeah thats what i was going to ask next, is there a particular case where you would look at a problem and go, "this problem SCREAMS for Log diff."

myininaya (myininaya):

\[y=sinx^{cosx}\] how would you find y' here then?

OpenStudy (anonymous):

take ln of both sides?

myininaya (myininaya):

this one screams it badly

OpenStudy (anonymous):

oh yeah. i would take ln of both sides. I guess ive just never called it by its name lol.

myininaya (myininaya):

\[\ln(y)=\ln(sinx)^{cosx}\] \[\ln(y)=cosx*\ln(sinx)\] \[\frac{y'}{y}=-sinx*\ln(sinx)+cosx*\frac{cosx}{sinx}\]

OpenStudy (anonymous):

what would the final anser be for the original problem i posted

myininaya (myininaya):

it would be what i have above multiplied by y (y=tan(theta)sqrt{2(theta)+1}

myininaya (myininaya):

lol

OpenStudy (anonymous):

wait what you orginal posted of the y=sinx^cosx

myininaya (myininaya):

now i was talking about your problem not the one i was doing for example

OpenStudy (anonymous):

i am confused, i still dont see what the final answer would be

myininaya (myininaya):

do you now see what i did for your problem?

myininaya (myininaya):

not*

OpenStudy (anonymous):

yea, above, before you did the example problem of y=sinx^cosx right?

myininaya (myininaya):

yes

myininaya (myininaya):

i got to y'/y=something well y' would just be that (something)*y

OpenStudy (anonymous):

so, you are saying multiply all of that by the orginal problem

myininaya (myininaya):

ok above when i was doing your problem i got y'/y=something right?

OpenStudy (anonymous):

right, cause we are trying to find the derivative so, we multiply both sides by y

myininaya (myininaya):

right!

myininaya (myininaya):

so we would then have y'=(something)*y right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

help with one more??

myininaya (myininaya):

make another post this one is getting long k? :)